# Second Order Stein-Variational Dynamic Optimization

### v1

@@ 5 Conclusion
In this paper, we have derived SVDDP based on the MEDDP formulation and SV Newton’s method, which has superior exploration capability to the original MEDDPs. Moreover, we incorporate state and control constraints with the barrier method to achieve constrained optimization with SVDDP. We compare SVDDP with a constrained version of MEDDP and MPPIs in three different systems. The results suggest that SVDDP can outperform MPPI in finding optimal solutions and in generating smooth state and control trajectories. Furthermore, SVDDP shows better exploration capability compared to MEDDPs. Future work includes hardware implementation with multimodal cost to validate the algorithm. Although we used the barrier function to represent constraints in this work, alternative optimization techniques can also be incorporated into SVDDP. In addition, while SVDDP in its current form requires differentiability, extensions that handle non-smooth dynamics [34] and non-smooth cost functions [35] are straightforward. Finally, theoretical analysis of the disruptive mechanism in Messy Dynamic Optimization and SVDDP is of great importance.

### v2

@@ 5 Conclusion
In this paper, we have derived SVDDP based on the MEDDP formulation and SVNM, which has superior exploration capability to the original MEDDPs. Moreover, we incorporate state and control constraints with the barrier method to achieve constrained optimization with SVDDP. We compare SVDDP with a constrained version of MEDDP in TO mode in three different systems, and also with MPPIs in two different systems in MPC mode. The results suggest that SVDDP can outperform MPPI in finding optimal solutions and in generating smooth state and control trajectories. Furthermore, SVDDP shows better exploration capability compared to MEDDPs. Future work includes hardware implementation with multimodal cost to validate the algorithm. Although we used the barrier function to represent constraints in this work, alternative optimization techniques can also be incorporated into SVDDP. In addition, while SVDDP in its current form requires differentiability, extensions that handle non-smooth dynamics [34] and non-smooth cost functions [35] are straightforward. Finally, theoretical analysis of the disruptive mechanism in Messy Dynamic Optimization and SVDDP is of great importance.